5 times the 5th term is 8 times the 8th term find its 12th term

Grade:

1 Answers

Sudheesh Singanamalla
13 years ago

5th term of AP = a+4d

8th term of AP = a+7d

5* (a+4d) = 8 * (a+7d)

5a + 20d = 8a + 56d

-3a = 36d

d = -a/12

12th term = a+ 11d

==> a + 11* -a / 12

==> ( 12a - 11a )/ 12

==> a/12

 

therfore the 12th term is a/12